OpenStudy (anonymous):

The probability of a student scoring 75% in class work is 0.64, and the probability of a student scoring 85% is 0.45. Event A: The student scores 75%. Event B: The student scores 85%. The probability of a student scoring 85% in class work, given that they have already scored 75% in class work, is 0.55. The probability of a student scoring 75% in class work, given that they have already scored 85% in class work, is 1. Which statement is true?

OpenStudy (hoblos):

if a student scored 85% it is sure that he scored 75% too so the second statement is absolutely true

OpenStudy (anonymous):

Events A and B are independent because P(A|B) = P(A). Events A and B are independent because P(A|B) = P(B). Events A and B are independent because P(A|B) = P(A) + P(B). Events A and B are not independent because P(A| B) ≠ P(A). Events A and B are not independent because P(A|B) = P(A).

OpenStudy (hoblos):

two events are independent if P(A|B) = P(A) here P(A|B) = 1 ≠ P(A) then Events A and B are not independent

OpenStudy (anonymous):

i still dont understand how to solve this.

OpenStudy (hoblos):

the first part is logical.. logically if it is given that a student scored 85% then he absolutely score more than 75% thus P(A|B) = 1 for the second part you have to find if they are independent or not two events A & B are independent if P(A|B) = P(A) P(A|B) = 1 P(A) = 0.64 thus P(A| B) ≠ P(A) and A & B are not independent so Events A and B are not independent because P(A|B) ≠ P(A)

OpenStudy (anonymous):

i see, thanks.

OpenStudy (hoblos):

welcome :)

OpenStudy (anonymous):

so what was the answer? @yomylola @hoblos

OpenStudy (anonymous):

D